# UniEmoX: Cross-modal Semantic-Guided Large-Scale Pretraining for Universal Scene Emotion Perception

### v2

@@ V-E Visualization
TABLE VII: The evaluation of various multimodal large models in simplified dialogue scenarios examines emotion classification results, causal analysis, and the performance of our proposed UniEmoX model. The table displays the results related to the left image in Fig. 10.
@@
TABLE VIII: The evaluation of various multimodal large models in simplified dialogue scenarios examines emotion classification results, causal analysis, and the performance of our proposed UniEmoX model. The table displays the results related to the right image in Fig. 10.
@@ V-F Comparative Analysis of Emotion Understanding Capabilities in Multimodal Large Models
    Multimodal large models demonstrate exceptional accuracy in emotion recognition and depth of emotional semantic understanding on most test samples. However, they still have limitations. For instance, some models demonstrate limited depth in emotional understanding. As shown in the left image of Fig. 10 and TABLE VII, most viewers might initially interpret the character’s facial expression as indicating happiness. Consequently, models such as ChatGLM, Baichuan, and SenseChat classify it as ’Joy’. However, this interpretation is superficial. A more nuanced analysis reveals that the character’s upward gaze suggests a reaction to something unexpected above. Therefore, the most accurate emotional label for this image is ’Surprise’. Notably, ERNIE Bot, iFlytek Spark, Qwen, and GPT-4o correctly identified this emotion.
@@
    As shown in the right image of Fig. 10 and TABLE VIII, the initial impression might be that the chaotic environment evokes a sense of ’disgust’, which aligns with the assessments of iFlytek Spark, ChatGLM, Qwen, Baichuan, and SenseChat. However, the presence of a crucial bloodstain and the eerie ambiance of the environment more accurately elicit a feeling of ’fear’. ERNIE Bot and GPT-4o correctly identified this emotion.
@@ VII Limitation
In constructing the emotional pretraining framework for visual perception, guided by three types of semantic information (namely, structural information in image space, textual-image semantic information, and psychological theory prior knowledge), we successfully develop an emotional pretraining framework tailored for large-scale complex environments (UniEmoX). While the experiments validate the progress of our UniEmoX, opportunities for enhancement remain. From the perspective of information granularity, the three types of semantic information in our UniEmoX are relatively coarse. Future efforts will focus on extracting finer-grained emotional-relevant regions from entire scene information, such as expressions, gazes, gaits, etc., to enhance the emotional expression capability of pretraining models. Additionally, future efforts will focus on constructing larger-scale and more diverse emotional datasets with richer emotional annotations.

### v3

@@ V-E Visualization
TABLE VIII: The evaluation of various multimodal large models in simplified dialogue scenarios examines emotion classification results, causal analysis, and the performance of our proposed UniEmoX model. The table displays the results related to the left image in Fig. 10.
@@
TABLE IX: The evaluation of various multimodal large models in simplified dialogue scenarios examines emotion classification results, causal analysis, and the performance of our proposed UniEmoX model. The table displays the results related to the right image in Fig. 10.
@@ V-F Comparative Analysis of Emotion Understanding Capabilities in Multimodal Large Models
    Multimodal large models demonstrate exceptional accuracy in emotion recognition and depth of emotional semantic understanding on most test samples. However, they still have limitations. For instance, some models demonstrate limited depth in emotional understanding. As shown in the left image of Fig. 10 and TABLE VIII, most viewers might initially interpret the character’s facial expression as indicating happiness. Consequently, models such as ChatGLM, Baichuan, and SenseChat classify it as ’Joy’. However, this interpretation is superficial. A more nuanced analysis reveals that the character’s upward gaze suggests a reaction to something unexpected above. Therefore, the most accurate emotional label for this image is ’Surprise’. Notably, ERNIE Bot, iFlytek Spark, Qwen, and GPT-4o correctly identified this emotion.
@@
    As shown in the right image of Fig. 10 and TABLE IX, the initial impression might be that the chaotic environment evokes a sense of ’disgust’, which aligns with the assessments of iFlytek Spark, ChatGLM, Qwen, Baichuan, and SenseChat. However, the presence of a crucial bloodstain and the eerie ambiance of the environment more accurately elicit a feeling of ’fear’. ERNIE Bot and GPT-4o correctly identified this emotion.
@@ VII Limitation
In constructing the emotional pretraining framework for visual perception, guided by three types of semantic information (namely, structural information in image space, textual-image semantic information, and psychological theory prior knowledge), we successfully develop an emotional pretraining framework tailored for large-scale complex environments. While the experiments validate the progress of our UniEmoX, opportunities for enhancement remain. From the perspective of information granularity, the three types of semantic information in our UniEmoX are relatively coarse. Future efforts will focus on extracting finer-grained emotional-relevant regions from entire scene information, such as expressions, gazes, gaits, etc., to enhance the emotional expression capability of pretraining models. Additionally, future efforts will focus on constructing larger-scale and more diverse emotional datasets with richer emotional annotations.
VIII Acknowledgments
This work was supported by Special Project of the National Natural Science Foundation of China (62441614), Anhui Province Key R&D Program (202304a05020068) and General Programmer of the National Natural Science Foundation of China (62376084).